i tried to work this out cna you please check
x^3+ax+3a+3x^2
i re order to get the like terms together this is what i came up with
3x^2+x^3+ax+3x^2
x(3x+x^2)(a+3a)
x(3+x)a(3)
i think i am wrong can you please show me an eash way or lend any tips you may have.

Answer by stanbon(75887)

x^3+ax+3a+3x^2
Put the x-terms in decending order as follows:
=x^3+3x^2+ax+3a
=x^2(x+3)+a(x+3)
=(x+3)(x^2+a)
cheers,
stan H.